# Benchmarking Humans and Machines on Complex Multilingual Speech Understanding Tasks

### v1

@@ 4.1 Speech stimuli
In our study, we investigate long-context speech understanding in humans and machines. To this end, we developed our own speech stimuli as the requisite data, in the form of read speech. We recorded a dataset featuring $20$ speakers, comprising $10$ native Kannada speakers and $10$ native Hindi speakers, balanced across age and gender to ensure diversity and representativeness. Both these sets of speakers were also multilingual with English as the secondary language. Since open source data of this nature in multi-lingual settings was not available, we resorted to record audio data for this study. Each speaker narrated $20$ stories: $10$ in English, and $10$ in their native language (Kannada or Hindi), spoken with prosody and fictional style. Each story, spanning approximately 400400450450 words, lasted 2.52.53.53.5 minutes in audio form, yielding a total of $20$ hours of audio data across the three languages. Our design of long-form story telling mimics real-world scenarios where sustained attention to longer contexts in spoken form is part of natural day-to-day human interactions. Each story is paired with $10$ manually constructed, factual and context-based multi-choice questions, each with four options. Post-recording, all audio clips were carefully trimmed to remove unwanted silence and other artifacts. Silences in the recordings were capped at $150$ milliseconds to maintain the natural flow of speech while minimizing disruptions that may impact listener attention or model performance in complex multi-speaker scenarios. Audio normalization was applied to achieve a consistent loudness level across all recordings. For mixed-channel trials, paired recordings were aligned to have approximately equal duration and energy balanced to maintain an average signal-to-interference ratio (SIR) of 0 dB for 5050ms segments. The two/three channel mixtures were obtained by adding mono-channel audio recordings. Further, all mixtures included recordings from both male and female speakers. This design allowed the inclusion of gender-based textual prompts to guide human or machine attention.
@@
Focus prompts: Prompt: Please focus only on the {MALE/FEMALE} speaker in the audio mixture and answer the following questions.
List of Questions: <Question-1,Options>, <Question-2,Options>, ... Instructions: Please follow the Focus prompt. Option: letter (A, B, C, D) corresponding to the selected answer. Evidence: one-line quote from the audio supporting the answer. Answer format: Question <index> Option: <A/B/C/D> Evidence: <one-line quote> Input Template: <Focus prompts>, <List of questions>, <Instructions>, <Answer format>, <Audio-data> Example Response: Question-1 Option:B Evidence: "the boy said he was late" Question-2 Option:C Evidence: "the traffic on the route was heavy."
@@ 4.2 Human Evaluation
During each behavioral session, participants were presented with mono and mixed-channel audio clips drawn from their native-language and English. We ensured that no listener was exposed to the same speaker or the same content more than once, throughout the listening period, thereby avoiding any familiarity bias.
For mono recordings, participants were instructed to listen to the story and answer a corresponding questionnaire. For mixed-channel recordings, participants were prompted to attend to the speech from a specified gender and answer the questionnaire. Each mixed-channel audio had $20$ questions in the questionnaire in a random order, with $10$ questions each from two of the audio recordings in the mixture (one male and one female speaker). Prior to the main sessions, participants completed a priming phase with a sample mono and a two-channel mixture with a practice questionnaires to familiarize them with the task. We record attended and unattended comprehension accuracy for the subsequent behavioral analysis.
@@ 4.3 Machine Evaluation
Table 2: Comparison of human and model evaluation results (Accuracy (%)). The human study involved $40$ participants with each participant listening to 22 mono and $2$ mixed channel trials per language, with each trial of duration 33 minutes. The model study involved $40$ trials in each language, consisting of 2020 mono and $20$ mixed-channel trials. Further, each trial had $10$ questions. Thus, the number of trials used in evaluation is identical for both humans and machines.
@@ 5 Results
The benchmarking results in Table 2 indicate that, with only minor exceptions, models achieve performance comparable to or exceed humans under mono-speaker conditions. The following are the key takeaways from the results presented in Table 2:

### v2

@@ 4.1 Speech stimuli
In our study, we investigate long-context speech understanding in humans and machines. To this end, we developed our own speech stimuli as the requisite data, in the form of read speech. We recorded a dataset featuring $20$ speakers, comprising $10$ native Kannada speakers and $10$ native Hindi speakers, balanced across age and gender to ensure diversity and representativeness. Both these sets of speakers were also multilingual with English as the secondary language. Since open source data of this nature in multi-lingual settings was not available, we resorted to record audio data for this study. Each speaker narrated $20$ stories: $10$ in English, and $10$ in their native language (Kannada or Hindi), spoken with prosody and fictional style. Each story, spanning approximately 400400450450 words, lasted 2.52.53.53.5 minutes in audio form, yielding a total of $20$ hours of audio data across the three languages. Our design of long-form story telling mimics real-world scenarios where sustained attention to longer contexts in spoken form is part of natural day-to-day human interactions. Each story is paired with $10$ manually constructed, factual and context-based multi-choice questions, each with four options. Post-recording, all audio clips were carefully trimmed to remove unwanted silence and other artifacts. Silences in the recordings were capped at $150$ milliseconds to maintain the natural flow of speech while minimizing disruptions that may impact listener attention or model performance in complex multi-speaker scenarios. Audio normalization was applied to achieve a consistent loudness level across all recordings. For mixed-channel trials, paired recordings were aligned to have approximately equal duration and energy balanced to maintain an average signal-to-interference ratio (SIR) of 0 dB for 5050ms segments, while no explicit control was imposed on gender- or speaker-specific acoustic factors. The two/three channel mixtures were obtained by adding mono-channel audio recordings. Further, all mixtures included recordings from both male and female speakers. This design allowed the inclusion of gender-based textual prompts to guide human or machine attention.
@@
Focus prompt: Prompt: Please focus only on the {MALE/FEMALE} speaker in the audio mixture and answer the following questions.
List of Questions: <Question-1,Options>, <Question-2,Options>, ... Instructions: Please follow the Focus prompt. Option: letter (A, B, C, D) corresponding to the selected answer. Evidence: one-line quote from the audio supporting the answer. Answer format: Question <index> Option: <A/B/C/D> Evidence: <one-line quote> Input Template: <Focus prompt>, <List of questions>, <Instructions>, <Answer format>, <Audio-data> Example Response: Question-1 Option:B Evidence: "the boy said he was late" Question-2 Option:C Evidence: "the traffic on the route was heavy."
@@ 4.2 Human Evaluation
During each session, participants were presented with mono and mixed-channel audio clips drawn from their native-language and English. We ensured that no listener was exposed to the same speaker or the same content more than once, throughout the listening period, thereby avoiding any familiarity bias.
Participants listened to mono recordings or attended to a prompted gender in mixed-channel recordings before answering the questionnaire. Each mixed-channel audio had $20$ questions in the questionnaire in a random order, with $10$ questions each from two of the audio recordings in the mixture (one male and one female speaker). Prior to the main sessions, participants completed a priming phase with a sample mono and a two-channel mixture with a practice questionnaires to familiarize them with the task. We record attended and unattended comprehension accuracy for the subsequent behavioral analysis.
@@ 4.3 Machine Evaluation
Table 2: Comparison of human and model evaluation results (Accuracy (%)). The human study involved $40$ participants (2020 Hindi-native, 2020 Kannada-native) with each participant listening to $1$ mono and $1$ mixed channel trials per language, with each trial of duration 33 minutes. The model study involved $40$ trials in each language, consisting of 2020 mono and $20$ mixed-channel trials. Further, each audio in a trial had $10$ questions. Thus, the number of trials used in evaluation is identical for both humans and machines. The color coding differentiates mono, mixture, attended, and unattended performance for both humans and models.
@@ 5 Results
The following are the key takeaways from the benchmarking results presented in Table 2: